OpenStudy (anonymous):

PRE-CAL; use the given pair of functions to find and simplify expressions for the following functions and state the domain of each using interval notation. • (gf)(x) • (f g)(x) • (f f)(x) 16. f(x)=3x-

OpenStudy (phoenixfire):

(gf)(x) means to multiply each equations (gf)(x)=g(x)*f(x) substitute the functions in and attempt to simplify.

OpenStudy (anonymous):

@PhoenixFire on the first one for (fg)(x)= 3√x-5 what would be my domain?

OpenStudy (anonymous):

@PhoenixFire (gf)(x)=√3x-5

OpenStudy (anonymous):

(ff)(x)=3(3x-5)-5 9x-15-5 9x-20

OpenStudy (phoenixfire):

no you got it wrong... it's not g(f(x)). the way you posed the question is (g*f)(x) f(x)*g(x)=(3x-5)(√x)=3x√x-5√x I don't actually know how to simplify that. but the domain is all positive values of X because √ can't have negatives